OpenStudy (anonymous):

Suppose the probability that it rains in the next two days are 1/3 for tomorrow and 1/6 for the day after tomorrow. What is p( rain tomorrow then rain the day after tomorrow ? 1/2 1/18 2/9 1/9 Show your work

OpenStudy (kropot72):

The required probability is the product of the two values of probability that are given.

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

\[P(rains\ both\ days)=\frac{1}{3}\times \frac{1}{6}=\ ?\]

OpenStudy (anonymous):

2/18

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

\[\frac{1}{3}\times \frac{1}{6}=\frac{1\times 1}{3\times 6}=\ ?\]

OpenStudy (anonymous):

1/18

OpenStudy (kropot72):

1/18 is correct :)

OpenStudy (anonymous):

thank you :)

OpenStudy (kropot72):

You're welcome :)